OpenStudy (anonymous):

how do you solve x over 2 = 3x over 4 + 5

OpenStudy (anonymous):

it's x/2 = 3x/4 + 5 the plus is not under the fraction

OpenStudy (anonymous):

x/2= 3x/4 +5 2x/4=3x/4 +5 (multiply left side by 2 to get common denominator) 2x=3x+20 (multiply both sides by 4 to get rid of the denominator) 0=x+20 (subtracted 2x from 3x and left zero on one side)

OpenStudy (anonymous):

deja vu all over again